Molly is 4 times as old as Jill. The sum of their ages is 60. What are their ages?

Found 2 solutions by mananth, jim_thompson5910:
Answer by mananth(16946)

Let Jill's age be j
Molly's age will be 4j
M + J = 60
plug Molly's age
4J+J=60
5J=60
J= 12
Jolly's age = 12 years
You find Molly's age
M = 4j

Answer by jim_thompson5910(35256)


Answers:
Molly is 48 years old
Jill is 12 years old

============================================

Work Shown:
M = Molly's age
J = Jill's age
Molly is 4 times old as Jill, meaning
Molly's age = 4*(Jill's age)
M = 4*J

Their ages add to 60
M+J = 60
4J + J = 60 ... replace M with 4J
5J = 60
5J/5 = 60/5 ... divide both sides by 5
J = 12
Jill is 12 years old

M = 4*J
M = 4*12
M = 48
Molly is 48 years old

Check:
M+J = 60
48+12 = 60
60 = 60
The answers are confirmed.